OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. (3 points) Question 1 options: 1) y - 6 = -3x - 3 2) y - 6 = -3(x + 1) 3) y = -3x + 3 4) 3x + y = 3

OpenStudy (anonymous):

y-y1=m(x-x1) where m=slope (x1,y1)= point given

OpenStudy (anonymous):

Plug that information into the equation I gave you and simplify accordingly.

OpenStudy (anonymous):

we know m=-3 (x1,y1)=(-1, 6)

OpenStudy (anonymous):

plug that into \[y-y1=m(x-x1)\]

OpenStudy (anonymous):

you guys gave the answer away....why?

OpenStudy (anonymous):

What i was looking for....